We know that limit of (3n/(n+4)) when n goes to infinity is 3. If the definition of convergence of sequences, ε=10^(-5) then determine the cutoff index N.
1
Expert's answer
2020-12-21T18:01:54-0500
"\big|\dfrac{3n}{n+4}-3\big|<10^{-5} \text{ whenever}\ n>N"

"\big|\dfrac{3n-3n-12}{n+4}\big|<10^{-5} \text{ whenever}\ n>N"

"\big|\dfrac{-12}{n+4}\big|<10^{-5} \text{ whenever}\ n>N"

"\dfrac{12}{10^{-5}}<n+4\text{ whenever}\ n>N"

"n>1200000-4"

"N=1199996"